OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. (4, 7) and (0, 7)

OpenStudy (anonymous):

@Kommander_Kitten

OpenStudy (anonymous):

This is a horizontal line, so the slope is 0.

OpenStudy (amorfide):

to work out the equation of a line, we will start by finding the gradient of the line also known as the slope \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] where \[(4,7)=(x _{1},y _{1})\] and \[(0,7)=(x _{2},y _{2})\] now the standard equation of a line \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

0? Alright, thank you. How would I write it in standard form though? That's what I cannot figure out.

OpenStudy (anonymous):

Plug in what you know: y-7=0(x-4)

OpenStudy (anonymous):

I suck at standard form

OpenStudy (anonymous):

Simplify, and get it into standard form.

OpenStudy (amorfide):

standard form is y=mx+c where m is the slope and c is the intercept

OpenStudy (anonymous):

That's not standard form. y=mx+b is slope-intercept form.

OpenStudy (amorfide):

lol i remember now, you are the person who doesn't want me to help you sorry

OpenStudy (anonymous):

Lol it's not that. I've been b*tchy these past few days to everybody. I'm sorry for that

OpenStudy (amorfide):

oh it is fine lol, but both are the slope intercept form, it just depends what book or teacher you get, some use y=mx+a good luck with your other questions

OpenStudy (anonymous):

I do online school so I have to teach myself -_- Lol I wish they gave me books. I just have the Internet

OpenStudy (anonymous):

I don't understand how to get it into that form though. I have it all wrote down, like all my work. It's just not clicking in my head lol

OpenStudy (amorfide):

yeah self teaching is hard anyways to finish off your question use \[y-y _{1}=m(x-x _{1})\] substitute your values in where (4,7)=(x1,y1) so you get y-7=0(x-4) y-7=0 y=7

OpenStudy (anonymous):

Write it in slope-intercept form, and then change it to standard form= y=0x+7----->y+0x=7

OpenStudy (anonymous):

The standard form and answer is y+0x=7

OpenStudy (anonymous):

THAT'S WHAT I HAD! It just didn't look right to me lol

OpenStudy (anonymous):

I wish I could give you both a medal!

OpenStudy (amorfide):

give credit to him

OpenStudy (amorfide):

he deserved it

OpenStudy (anonymous):

Okay. But thank you, too, @amorfide !

OpenStudy (amorfide):

anytime just tag me whenever you need help